\begin{thebibliography}{26}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Bahdanau et~al.(2014)Bahdanau, Cho, and Bengio]{attention}
Dzmitry Bahdanau, Kyunghyun Cho, and Yoshua Bengio.
\newblock Neural machine translation by jointly learning to align and
  translate.
\newblock \emph{CoRR}, abs/1409.0473, 2014.
\newblock URL \url{http://arxiv.org/abs/1409.0473}.

\bibitem[Chen et~al.(2015)Chen, Wilson, Tyree, Weinberger, and Chen]{chen:2015}
Wenlin Chen, James Wilson, Stephen Tyree, Kilian Weinberger, and Yixin Chen.
\newblock Compressing neural networks with the hashing trick.
\newblock In \emph{International Conference on Machine Learning}, pages
  2285--2294, 2015.

\bibitem[Cheng et~al.(2015)Cheng, Yu, Feris, Kumar, Choudhary, and
  Chang]{yu:15}
Yu~Cheng, Felix~X. Yu, Rog{\'{e}}rio~Schmidt Feris, Sanjiv Kumar, Alok~N.
  Choudhary, and Shih{-}Fu Chang.
\newblock An exploration of parameter redundancy in deep networks with
  circulant projections.
\newblock In \emph{2015 {IEEE} International Conference on Computer Vision,
  {ICCV} 2015, Santiago, Chile, December 7-13, 2015}, pages 2857--2865, 2015.
\newblock \doi{10.1109/ICCV.2015.327}.
\newblock URL \url{https://doi.org/10.1109/ICCV.2015.327}.

\bibitem[Chu and Liu(1965)]{mst}
Chu and Liu.
\newblock On the shortest arborescence of a directed graph.
\newblock \emph{Science Sinica}, Vol.14, 1965.

\bibitem[Dozat and Manning(2016)]{deepbiaffine}
Timothy Dozat and Christopher~D. Manning.
\newblock Deep biaffine attention for neural dependency parsing.
\newblock \emph{CoRR}, abs/1611.01734, 2016.
\newblock URL \url{http://arxiv.org/abs/1611.01734}.

\bibitem[Dozat et~al.(2017)Dozat, Qi, and Manning]{dozat2017stanford}
Timothy Dozat, Peng Qi, and Christopher~D Manning.
\newblock Stanford's graph-based neural dependency parser at the conll 2017
  shared task.
\newblock \emph{Proceedings of the CoNLL 2017 Shared Task: Multilingual Parsing
  from Raw Text to Universal Dependencies}, pages 20--30, 2017.

\bibitem[Edmonds(1967)]{edmonds}
Jack Edmonds.
\newblock Optimum branchings.
\newblock \emph{JOURNAL OF RESEARCH of the National Bureau of Standards},
  71\penalty0 (4), 1967.

\bibitem[Gray et~al.(2006)]{gray2006toeplitz}
Robert~M Gray et~al.
\newblock Toeplitz and circulant matrices: A review.
\newblock \emph{Foundations and Trends{\textregistered} in Communications and
  Information Theory}, 2\penalty0 (3):\penalty0 155--239, 2006.

\bibitem[Hashimoto et~al.(2017)Hashimoto, xiong, Tsuruoka, and
  Socher]{bilinear-attention}
Kazuma Hashimoto, caiming xiong, Yoshimasa Tsuruoka, and Richard Socher.
\newblock A joint many-task model: Growing a neural network for multiple nlp
  tasks.
\newblock In \emph{Proceedings of the 2017 Conference on Empirical Methods in
  Natural Language Processing}, pages 1923--1933. Association for Computational
  Linguistics, 2017.
\newblock URL \url{http://aclweb.org/anthology/D17-1206}.

\bibitem[Hayashi and Shimbo(2017)]{hayashi:17-2}
Katsuhiko Hayashi and Masashi Shimbo.
\newblock On the equivalence of holographic and complex embeddings for link
  prediction.
\newblock In \emph{Proceedings of the 55th Annual Meeting of the Association
  for Computational Linguistics, {ACL} 2017, Vancouver, Canada, July 30 -
  August 4, Volume 2: Short Papers}, pages 554--559, 2017.
\newblock \doi{10.18653/v1/P17-2088}.
\newblock URL \url{https://doi.org/10.18653/v1/P17-2088}.

\bibitem[Hinton et~al.(2015)Hinton, Vinyals, and Dean]{hinton:15}
Geoffrey~E. Hinton, Oriol Vinyals, and Jeffrey Dean.
\newblock Distilling the knowledge in a neural network.
\newblock \emph{CoRR}, abs/1503.02531, 2015.

\bibitem[Hubara et~al.(2016)Hubara, Courbariaux, Soudry, El{-}Yaniv, and
  Bengio]{itay:16}
Itay Hubara, Matthieu Courbariaux, Daniel Soudry, Ran El{-}Yaniv, and Yoshua
  Bengio.
\newblock Quantized neural networks: Training neural networks with low
  precision weights and activations constrained to +1 or -1.
\newblock \emph{CoRR}, abs/1609.07061, 2016.
\newblock URL \url{http://arxiv.org/abs/1609.07061}.

\bibitem[Ishihara et~al.(2018)Ishihara, Hayashi, Manabe, Shimbo, and
  Nagata]{ishihara:18}
Takahiro Ishihara, Katsuhiko Hayashi, Hitoshi Manabe, Masashi Shimbo, and
  Masaaki Nagata.
\newblock Neural tensor networks with diagonal slice matrices.
\newblock In \emph{Proceedings of the 2018 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, {NAACL-HLT} 2018, New Orleans, Louisiana, USA, June 1-6, 2018,
  Volume 1 (Long Papers)}, pages 506--515, 2018.
\newblock URL \url{https://aclanthology.info/papers/N18-1047/n18-1047}.

\bibitem[Kiperwasser and Goldberg(2016)]{mlp-attention}
Eliyahu Kiperwasser and Yoav Goldberg.
\newblock Simple and accurate dependency parsing using bidirectional lstm
  feature representations.
\newblock \emph{Transactions of the Association for Computational Linguistics},
  4:\penalty0 313--327, 2016.
\newblock URL \url{http://aclweb.org/anthology/Q16-1023}.

\bibitem[Liu et~al.(2017)Liu, Wu, and Yang]{liu2017analogy}
Hanxiao Liu, Yuexin Wu, and Yiming Yang.
\newblock Analogical inference for multi-relational embeddings.
\newblock \emph{CoRR}, abs/1705.02426, 2017.
\newblock URL \url{http://arxiv.org/abs/1705.02426}.

\bibitem[Lu et~al.(2016)Lu, Sindhwani, and Sainath]{lu:16}
Zhiyun Lu, Vikas Sindhwani, and Tara~N Sainath.
\newblock Learning compact recurrent neural networks.
\newblock In \emph{Acoustics, Speech and Signal Processing (ICASSP), 2016 IEEE
  International Conference on}, pages 5960--5964. IEEE, 2016.

\bibitem[{Ma} et~al.(2018){Ma}, {Hu}, {Liu}, {Peng}, {Neubig}, and
  {Hovy}]{stackpointer}
X.~{Ma}, Z.~{Hu}, J.~{Liu}, N.~{Peng}, G.~{Neubig}, and E.~{Hovy}.
\newblock {Stack-Pointer Networks for Dependency Parsing}.
\newblock \emph{ArXiv e-prints}, May 2018.

\bibitem[Manabe et~al.(2018)Manabe, Hayashi, and
  Shimbo]{DBLP:conf/aaai/ManabeHS18}
Hitoshi Manabe, Katsuhiko Hayashi, and Masashi Shimbo.
\newblock Data-dependent learning of symmetric/antisymmetric relations for
  knowledge base completion.
\newblock In \emph{Proceedings of the Thirty-Second {AAAI} Conference on
  Artificial Intelligence, New Orleans, Louisiana, USA, February 2-7, 2018},
  2018.
\newblock URL
  \url{https://www.aaai.org/ocs/index.php/AAAI/AAAI18/paper/view/16211}.

\bibitem[Nickel et~al.(2011)Nickel, Tresp, and Kriegel]{rescal}
Maximilian Nickel, Volker Tresp, and Hans-Peter Kriegel.
\newblock A three-way model for collective learning on multi-relational data.
\newblock In \emph{ICML}, volume~11, pages 809--816, 2011.

\bibitem[Nickel et~al.(2015)Nickel, Rosasco, and Poggio]{hole}
Maximilian Nickel, Lorenzo Rosasco, and Tomaso~A. Poggio.
\newblock Holographic embeddings of knowledge graphs.
\newblock \emph{CoRR}, abs/1510.04935, 2015.
\newblock URL \url{http://arxiv.org/abs/1510.04935}.

\bibitem[Rush et~al.(2015)Rush, Chopra, and Weston]{attention-summarization}
Alexander~M Rush, Sumit Chopra, and Jason Weston.
\newblock A neural attention model for abstractive sentence summarization.
\newblock \emph{arXiv preprint arXiv:1509.00685}, 2015.

\bibitem[Sindhwani et~al.(2015)Sindhwani, Sainath, and Kumar]{vikas:15}
Vikas Sindhwani, Tara~N. Sainath, and Sanjiv Kumar.
\newblock Structured transforms for small-footprint deep learning.
\newblock In \emph{Advances in Neural Information Processing Systems 28: Annual
  Conference on Neural Information Processing Systems 2015, December 7-12,
  2015, Montreal, Quebec, Canada}, pages 3088--3096, 2015.

\bibitem[Socher et~al.(2013)Socher, Chen, Manning, and Ng]{ntn}
Richard Socher, Danqi Chen, Christopher~D. Manning, and Andrew~Y. Ng.
\newblock Reasoning with neural tensor networks for knowledge base completion.
\newblock In \emph{Advances in Neural Information Processing Systems 26: 27th
  Annual Conference on Neural Information Processing Systems 2013. Proceedings
  of a meeting held December 5-8, 2013, Lake Tahoe, Nevada, United States.},
  pages 926--934, 2013.

\bibitem[Trouillon et~al.(2016)Trouillon, Welbl, Riedel, Gaussier, and
  Bouchard]{Trouillon:2016:CES:3045390.3045609}
Th{\'e}o Trouillon, Johannes Welbl, Sebastian Riedel, \'{E}ric Gaussier, and
  Guillaume Bouchard.
\newblock Complex embeddings for simple link prediction.
\newblock In \emph{Proceedings of the 33rd International Conference on
  International Conference on Machine Learning - Volume 48}, ICML'16, pages
  2071--2080. JMLR.org, 2016.
\newblock URL \url{http://dl.acm.org/citation.cfm?id=3045390.3045609}.

\bibitem[Trouillon et~al.(2017)Trouillon, Dance, Gaussier, Welbl, Riedel, and
  Bouchard]{complex-journal}
Th{\'{e}}o Trouillon, Christopher~R. Dance, {\'{E}}ric Gaussier, Johannes
  Welbl, Sebastian Riedel, and Guillaume Bouchard.
\newblock Knowledge graph completion via complex tensor factorization.
\newblock \emph{Journal of Machine Learning Research}, 18:\penalty0
  130:1--130:38, 2017.
\newblock URL \url{http://jmlr.org/papers/v18/papers/v18/16-563.html}.

\bibitem[Zeman et~al.(2017)Zeman, Popel, Straka, Hajic, Nivre, Ginter,
  Luotolahti, Pyysalo, Petrov, Potthast, Tyers, Badmaeva, Gokirmak, Nedoluzhko,
  Cinkova, Hajic~jr., Hlavacova, Kettnerov{\'a}, Uresova, Kanerva, Ojala,
  Missil{\"a}, Manning, Schuster, Reddy, Taji, Habash, Leung, de~Marneffe,
  Sanguinetti, Simi, Kanayama, dePaiva, Droganova, Mart{\'i}nez~Alonso,
  {\c{C}}{\"o}ltekin, Sulubacak, Uszkoreit, Macketanz, Burchardt, Harris,
  Marheinecke, Rehm, Kayadelen, Attia, Elkahky, Yu, Pitler, Lertpradit, Mandl,
  Kirchner, Alcalde, Strnadov{\'a}, Banerjee, Manurung, Stella, Shimada, Kwak,
  Mendonca, Lando, Nitisaroj, and Li]{conll2017}
Daniel Zeman, Martin Popel, Milan Straka, Jan Hajic, Joakim Nivre, Filip
  Ginter, Juhani Luotolahti, Sampo Pyysalo, Slav Petrov, Martin Potthast,
  Francis Tyers, Elena Badmaeva, Memduh Gokirmak, Anna Nedoluzhko, Silvie
  Cinkova, Jan Hajic~jr., Jaroslava Hlavacova, V{\'a}clava Kettnerov{\'a},
  Zdenka Uresova, Jenna Kanerva, Stina Ojala, Anna Missil{\"a}, Christopher~D.
  Manning, Sebastian Schuster, Siva Reddy, Dima Taji, Nizar Habash, Herman
  Leung, Marie-Catherine de~Marneffe, Manuela Sanguinetti, Maria Simi, Hiroshi
  Kanayama, Valeria dePaiva, Kira Droganova, H{\'e}ctor Mart{\'i}nez~Alonso,
  {\c{C}}a{\u{g}}r{\i} {\c{C}}{\"o}ltekin, Umut Sulubacak, Hans Uszkoreit,
  Vivien Macketanz, Aljoscha Burchardt, Kim Harris, Katrin Marheinecke, Georg
  Rehm, Tolga Kayadelen, Mohammed Attia, Ali Elkahky, Zhuoran Yu, Emily Pitler,
  Saran Lertpradit, Michael Mandl, Jesse Kirchner, Hector~Fernandez Alcalde,
  Jana Strnadov{\'a}, Esha Banerjee, Ruli Manurung, Antonio Stella, Atsuko
  Shimada, Sookyoung Kwak, Gustavo Mendonca, Tatiana Lando, Rattima Nitisaroj,
  and Josie Li.
\newblock Conll 2017 shared task: Multilingual parsing from raw text to
  universal dependencies.
\newblock In \emph{Proceedings of the CoNLL 2017 Shared Task: Multilingual
  Parsing from Raw Text to Universal Dependencies}, pages 1--19. Association
  for Computational Linguistics, 2017.
\newblock \doi{10.18653/v1/K17-3001}.
\newblock URL \url{http://www.aclweb.org/anthology/K17-3001}.

\end{thebibliography}

\documentclass[11pt]{article}
\usepackage{paclic32}
\usepackage{times}
\usepackage{latexsym}
\usepackage{amsmath}
\usepackage{multirow}
\usepackage{url}
\usepackage[round]{natbib}

\usepackage{graphicx, xcolor}
\usepackage{latexsym}
\usepackage{booktabs}
\usepackage{amsmath,amssymb}
\usepackage{tikz}
\usepackage{multirow,bigdelim}
\newcommand{\FFT}{\mathop{\mathfrak{F}}}

\def\Underline{\setbox0\hbox\bgroup\let\\\endUnderline}
\def\endUnderline{\vphantom{y}\egroup\smash{\underline{\box0}}\\}
\def\|{\verb\|}

\newcommand{\Cset}{\mathbb{C}}
\newcommand{\Rset}{\mathbb{R}}
\newcommand{\mat}[1]{\boldsymbol{\mathbf{#1}}}

\newcommand{\transpose}{^{\mathrm{T}}}

\newcommand{\diag}{\mathop{\text{diag}}}

\title{Reduction of Parameter Redundancy in Biaffine Classifiers \\ with Symmetric and Circulant Weight Matrices}

\author{Tomoki Matsuno$^{1}$, Katsuhiko Hayashi$^{2, 4}$, Takahiro Ishihara$^{1}$, Hitoshi Manabe$^{3}$, Yuji Matsumoto$^{1, 4}$ \\
  $^{1}$Nara Institute of Science and Technology \\
  $^{2}$Osaka University  $^{3}$Works Applications \\
  $^4$RIKEN Center for Advanced Intelligence Project\\
  \tt $^{1}$\{matsuno.tomoki.mr1,ishihara.takahiro.in0, matsu\}@is.naist.jp\\
  \tt $^{2}$khayashi0201@gmail.com $^{3}$manabe\_h@worksap.co.jp\\
}

\date{}

\begin{document}
\maketitle
\begin{abstract}
  Currently, the biaffine classifier has been attracting attention as a method to introduce an attention mechanism into the modeling of binary relations. For instance, in the field of dependency parsing, the Deep Biaffine Parser by \citeauthor{deepbiaffine} has achieved state-of-the-art performance as a graph-based dependency parser on the English Penn Treebank and CoNLL 2017 shared task. On the other hand, it is reported that parameter redundancy in the weight matrix in biaffine classifiers, which has $O(n^2)$ parameters, results in overfitting ($n$ is the number of dimensions). In this paper, we attempted to reduce the parameter redundancy by assuming either symmetry or circularity of weight matrices. In our experiments on the CoNLL 2017 shared task dataset, our model achieved better or comparable accuracy on most of the treebanks with more than 16\% parameter reduction.
\end{abstract}

\section{Introduction}

  Recently, methods based on attention mechanisms have been used in various fields of natural language processing \citep{attention, attention-summarization}. In tasks which deal with \textbf{binary relations} that take two words as arguments, e.g., dependency parsing, a good number of these models have achieved high performance \citep{mlp-attention, bilinear-attention, deepbiaffine}.
  
  \textbf{Biaffine transformation} is a method to incorporate an attention mechanism into binary relations proposed by \cite{deepbiaffine} (following them, we call this method \textbf{biaffine classifier}). They achieved the state-of-the-art performance among graph-based dependency parsers for the English Penn Treebank. In addition, the state-of-the-art transition-based parser on English Penn Treebank uses a biaffine classifier to evaluate the probability distribution of a word coming into the stack at a step point~\citep{stackpointer}.
  
  While biaffine transformation has rich expressiveness in modeling binary relations, its number of parameters in the weight matrix (bilinear term) is $O(n^2)$ (where $n$ is the number of dimensions). This redundant parameters can result in high degree of freedom of the model, thus causing overfitting especially when a large number of training samples are not available~\citep{hole}.  
  
  In this paper, we attempt to reduce the redundancy by introducing the assumption of either symmetry or circularity in the weight matrix at a biaffine classifier. With either assumption, we can vectorize the matrix and reduce the space complexity to $O(n)$. Additionally, the time complexity becomes $O(n)$ in the case of symmetry and $O(n\log n)$ in the case of circularity with the fast Fourier transform. 
Furthermore, while the expressiveness of the model based on symmetry is restricted\footnote{Denote the score of the bilinear term by $f(a,b)$ when a word pair $a$, $b$ has a dependency relation. In this case, if the bilinear term has symmetry, $f(a,b)=f(b,a)$ . Therefore, this is not appropriate for expression of a directed edge. },  one based on circularity is able to express asymmetry relations.

In our experiments, we imposed constraints on the biaffine classifiers of the deep biaffine parser and examined the effect on the accuracy of the model. For our experiments, we used the dependency parsing dataset from the CoNLL 2017 shared task~\citep{conll2017}. We chose four languages which have relatively rich training examples and four languages which have fewer. From our experiments, we found that the method with the circularity assumption outperformed the baseline in most of the languages.

\section{Deep Biaffine Parser}

Models introduced in this paper are based on the \textbf{Deep Biaffine Parser} proposed by \citep{deepbiaffine}. They achieved the state-of-the-art accuracy on the CoNLL 2017 shared task for Universal Dependencies~\citep{conll2017}. 

This model receives a sequence of words and POS tags, and calculates the probability of an arc between each pair of words as well as a syntactic function label for each arc. For evaluation of scores, it uses Long Short-Term Memories (LSTMs), Multi-Layer Perceptrons (MLPs) and biaffine classifiers.

In the following sections, first, we explain the biaffine transformation which is the essential part of a biaffine classifier while skipping explanation about LSTM and MLP for the sake of simplicity. Then, we describe the overview of the model. It is worth noting that the structure of the model is different from that in \citep{deepbiaffine} in that it utilizes character level information. This is because we used an updated version of the model that was made for the shared task \citep{dozat2017stanford}. 

\subsection{Biaffine Transformation}

For the dependency parsing score functions, we use the biaffine transformation shown below to model binary relations. Here, $\oplus$ stands for concatenation of vectors. The first term on the right side represents relatedness score, and the second term the score of $v_i$ and $v_j$ appearing independently.  $b$ is bias term.

\begin{equation}\label{eq:biaffine}
\begin{split}
g(\mathbf{v}_i, \mathbf{v}_j) 
&= 
\mathbf{v}_i\transpose\mathbf{Av}_j + (\mathbf{v}_i \oplus \mathbf{v}_j) \transpose \mathbf{b} + b
\end{split}
\end{equation}

\subsection{Structure of the Model}

We show the structure of this model below.

\begin{enumerate}

\item
First, this model takes two sequences: words and POS tags. It uses a unidirectional LSTM to encode each words' character-level information into a vector. It then sums this vector with a separate token-level word embeddings. 

 \item
It then concatenates the vectors obtained above with POS embeddings and encodes them via a three layer bidirectional LSTM. $y_i$ denotes a vector made by concatenating the hidden states (not including cell states) from the LSTMs of both directions which corresponds to the $i$ th word $w_i$.

\item
These outputs are then transformed with MLPs. Here, we use distinct MLPs for dependents and heads in the prediction of arcs and labels.
\begin{equation}
\begin{split}
\mathbf{v}_{i}^{arc-head} &= \mathrm{MLP}^{arc-head}(\mathbf{y}_i),\\
\mathbf{v}_j^{arc-dep} &= \mathrm{MLP}^{arc-dep}(\mathbf{y}_j), \\
\mathbf{v}_{i}^{label-head} &= \mathrm{MLP}^{label-head}(\mathbf{y}_i),\\
\mathbf{v}_j^{label-dep} &= \mathrm{MLP}^{label-dep}(\mathbf{y}_j).
\end{split}
\end{equation}

where $\mathbf{v}_{i}^{arc-head}, \mathbf{v}_{j}^{arc-dep} \in \mathbb{R}^{n}$ \\
and $\mathbf{v}_{i}^{label-head}, \mathbf{v}_{j}^{label-dep} \in \mathbb{R}^{m}$ \\

\item
The scores of arcs between each word pair are calculated using a biaffine transformation.
\begin{equation} \label{eq:arc}
s_{i,j}^{(arc)} = {\mathbf{v}_i^{arc-head}}^{\mathrm T} \mathbf W \mathbf{v}_j^{arc-dep} + {\mathbf{v}_i^{arc-head}}^{\mathrm T} \mathbf b^{(arc)}.
\end{equation}

\item
Running Chu-Liu/Edmonds algorithm \citep{mst, edmonds} on the scores calculated in (3), we obtain a tree structure that maximizes the total score.

\item
It evaluates a score $s_{i,j}^{(l)}$  of assigning a label $l$ $(l \in \{1, 2, ..., L\};$ $L$: the number of labels) on the arc between the dependent word $w_i$ and its predicted head word $w_j$. The equation is defined below.

\begin{equation} \label{eq:label}
\begin{split}
s_{i,j}^{(l)} 
&= {\mathbf{v}_{i}^{label-head}}\transpose \mathbf{U}^{[l]} {\mathbf v_j^{label-dep}} \\
&+ {({\mathbf{v}_{i}^{label-head}} \oplus \mathbf v_j^{label-dep})}\transpose \mathbf{b}^{[l]} \\
&+ b^{[l]}.
\end{split}
\end{equation}

Here, a distinct weight matrix $\mathbf{U}^{[l]}$, weight vector $\mathbf{b}^{[l]}$ and bias $b^{[l]}$ are used for each label. The first term on the right side of the equation (4) represents the score of assigning the label $l$ to the arc between dependent $w_i$ and head $w_j$. The second term expresses the score of the label when the dependent and head are given independently.

\end{enumerate}

In our experiment, $\mat{W}\in\Rset^{n\times n}$ and $\mat{U}^{[l]}\in\Rset^{m\times m}$ account for about 17\% of the parameters in the deep biaffine parser. By reducing these parameters, we can expect not only improved memory efficiency but also less overfitting.

\section{Proposed Methods}
\label{sec:proposed}
In this section, we introduce our two proposed methods to reduce the number of parameters of the model. We impose either a symmetry or circularity constraint on the weight matrix $\mathbf{W}$ of (\ref{eq:arc}) and $\mathbf{U}^{[l]}~(\forall l\in\{1,2,\dots,L\})$ of (\ref{eq:label}).

\subsection{Symmetric Matrix Constraint}
This method assumes that the weight matrices of the bilinear terms are \textbf{symmetric matrices} and thus are diagonalizable. As a result, we can transform the bilinear term of the score functions into a ``triple inner product'' of two input vectors and a weight vector.
\subsubsection{Diagonalization of the weight matrix in the bilinear term}
When a matrix $\mathbf{W} \in \mathbb{R}^{n \times n}$ is symmetric, it can be diagonalized by an orthogonal matrix $\mathbf{O} \in \mathbb{R}^{n \times n}$ as below:
\[
\mathbf{W} = \mathbf{O} {\diag(\mathbf{w})} \mathbf{O}\transpose
\]
where $\mathbf{w} \in \mathbb{R}^{n}$ consists of the eigenvalues of $\mathbf{W}$ and $\diag(\mathbf{w}) \in \mathbb{R}^{n \times n}$ represents the diagonal matrix whose diagonal elements are $\mathbf{w}$. With this property, we can rewrite the bilinear term as follows:
\begin{equation}
\begin{split}
\label{eq:diag}
\mathbf{v}_i\transpose \mathbf{W} \mathbf{v}_j
&=
\mathbf{v}_i\transpose \mathbf{O} \diag(\mathbf{w}) \mathbf{O} \transpose \mathbf{v}_j  \\ &=
{\mathbf{v}^{\prime}_{i}}\transpose \diag(\mathbf{w}) \mathbf{v}^{\prime}_{j} \\
&=
\langle\mathbf{v}^{\prime}_{i}, \mathbf{w}, \mathbf{v}^{\prime}_{j}\rangle.
\end{split}
\end{equation}
where, $\mathbf{v}^{\prime}_{i}=\mathbf{O}\transpose \mathbf{v}_i$ and $\mathbf{v}^{\prime}_{j} = \mathbf{O}\transpose \mathbf{v}_j$, assuming $\mat{O}$ is learned implicitly. 
$\langle\mathbf{v}^{\prime}_{i}, \mathbf{w}, \mathbf{v}^{\prime}_{j}\rangle$ is a ``triple inner product'' of $\mathbf{v}^{\prime}_{i}, \mathbf{w}$ and $\mathbf{v}^{\prime}_{j}$ defined by $\langle \mathbf{a}, \mathbf{b}, \mathbf{c} \rangle = \sum_{k=1}^n a_k b_k c_k$. Consequently, the symmetry constraint on the matrix can reduce the number of weight parameters from $n^2$ to $n$.

\subsubsection{Simultaneous Diagonalization}
\label{sim-diag}
When a set of symmetric matrices forms a commuting family, they can be diagonalized by the same orthogonal matrix~\citep{liu2017analogy}. 
So we assume the weight matrices $\mathbf{U}^{[1]}, \mathbf{U}^{[2]}, \dots, \mathbf{U}^{[L]}$ of the scoring functions (\ref{eq:label}) form a commuting family. Namely, we assume:
\[
\mathbf{U}^{[p]}\mathbf{U}^{[q]} = \mathbf{U}^{[q]}\mathbf{U}^{[p]},~\forall p, q \in \{1,2, \dots, L\}.
\]
With this assumption, all $L$ weight matrices can be diagonalized simultaneously. Therefore, the vectors $\mathbf{v}_{i}^{label-head}$ and $\mathbf{v}_{j}^{label-dep}$ can be mapped by the same orthogonal matrix for all score functions.
\subsubsection{Score Functions}
Based on the above, under the assumption that all weight matrices are symmetric, we substitute the bilinear term in the biaffine transformations with a triple inner product. First, the score function for arc is defined as follows:
\begin{equation}
\label{eq:1arc}
\begin{split}
s_{i,j}^{(arc)} &= \langle {\mathbf{v}_i^{arc-head}}, \mathbf{w}, \mathbf{v}_j^{arc-dep} \rangle \\
&+ ({\mathbf v_{i}^{arc-head} \oplus \mathbf v_j^{arc-dep}})^{\mathrm T} \mathbf{b}
.
\end{split}
\end{equation}
where, $\mathbf{w} \in \mathbb{R}^n,~\mathbf{b} \in \mathbb{R}^{2n}$. Unlike in (\ref{eq:arc}), the second term of this function contains the arc-dep vector because we confirm that it improves the performance.

Scoring functions for labels are defined as follows:
\begin{equation} \label{eq:1label}
\begin{split}
s_{i,j}^{(l)} &= \langle {\mathbf v_{i}^{label-head}}, \mathbf{u}^{[l]}, {\mathbf v_j^{label-dep}} \rangle  \\
&+ ({\mathbf v_{i}^{label-head} \oplus \mathbf v_j^{label-dep}})^{\mathrm T} \mathbf {b}^{[l]}
.
\end{split}
\end{equation}
where, $\mathbf{u}^{[l]} \in \mathbb{R}^{m}~,\mathbf{b}^{[l]} \in \mathbb{R}^{2m}$. We eliminate the bias term $b^{[l]}$ in (\ref{eq:label}) because we confirm that it does not affect the performance.

\subsection{Circulant Matrix Constraint}
\cite{hole} used a circulant matrix for the bilinear transformation in knowledge graph completion model~\citep{rescal} to reduce the number of parameters and improve the computational efficiency. Following this method, we assume the weight matrices of the bilinear term in the biaffine transformations are circulant and propose new scoring functions based on that.

\subsubsection{Bilinear Transformation Using a Circulant Matrix}
We define the circulant matrix $C(\mathbf{w}) \in \mathbb{R}^{n \times n}$ for a vector $\mathbf{w} \in \mathbb{R}^{n} $ as follows:
\begin{equation}
C(\mathbf{w})=
\begin{bmatrix}
    w_1       & w_n & \dots & w_3 & w_2 \\
    w_2       & w_1 & w_n &  & w_3 \\
    \vdots    & w_2 & w_1 & \ddots & \vdots \\
    w_{n-1}   &     & \ddots & \ddots & w_{n} \\
    w_{n}     &  w_{n-1} & \dots & w_{2} & w_{1}
\end{bmatrix}
.
\end{equation}
where, $\mathbf{w} \transpose = (w_1, \dots, w_n)$. . Then, we replace the bilinear term with one where the weight matrix is a circulant matrix $C(w)$ with $n$ parameters:

\begin{equation}
\label{eq:circ-bilinear}
\mathbf{v}_{i} \transpose C(\mathbf{w})  \mathbf{v}_{j}.
\end{equation}

\subsubsection{Score Functions}
We propose score functions that employ (\ref{eq:circ-bilinear}) as the bilinear term in the biaffine transformation. The score function for an arc is then defined as follows:

\begin{equation}
\label{eq:2arc}
\begin{split}
s_{i,j}^{(arc)} &= {\mathbf{v}_i^{arc-head}} C(\mathbf{w}) \mathbf{v}_j^{arc-dep}  \\
&+ ({\mathbf v_{i}^{arc-head} \oplus \mathbf v_j^{arc-dep}})^{\mathrm T} \mathbf{b}
.
\end{split}
\end{equation}
where, $\mathbf{w} \in \mathbb{R}^n,~\mathbf{b} \in \mathbb{R}^{2n}$.

The score functions for labels are defined as follows:
\begin{equation} \label{eq:2label}
\begin{split}
s_{i,j}^{(l)} &= {\mathbf v_{i}^{label-head}}C(\mathbf{u}^{[l]}){\mathbf v_j^{label-dep}} \\
&+ ({\mathbf v_{i}^{label-head} \oplus \mathbf v_j^{label-dep}})\transpose \mathbf{b}^{[l]}
.
\end{split}
\end{equation}
where, $\mathbf{u}^{[l]} \in \mathbb{R}^{m}~,\mathbf{b}^{[l]} \in \mathbb{R}^{2m}$.

\subsubsection{Efficient Computation Using Fast Fourier Transformation}
In this section, we explain how to compute (\ref{eq:circ-bilinear}) efficiently using a fast Fourier transformation (FFT). We denote an $n$-point discrete Fourier transformation (DFT) matrix as $\FFT _n\in \mathbb{C}^{n \times n}$. Then, any circulant matrix $C(\mathbf{w}) \in \mathbb{R}^{n \times n}$ can be diagonalized as follows~\citep{gray2006toeplitz}:

\[
C(\mathbf{w}) = {\FFT}_n^{-1}\diag({\FFT}_n \mathbf{w}){\FFT}_n.
\]

With this property, we can rewrite (\ref{eq:circ-bilinear}) as a triple hermitian inner product ~\citep{liu2017analogy}:
\begin{align}
\label{eq:prop2}
\mathbf{v}_{i} \transpose C(\mathbf{w}) \mathbf{v}_{j}
&= \mathbf{v}_{i} {\FFT}_n^{-1}\diag({\FFT}_n \mathbf{w}){\FFT}_n \mathbf{v}_{j} \nonumber\\
&= \frac{1}{n} {\overline{{\FFT}_n \mathbf{v}_{i}}}\transpose \diag({\FFT}_n \mathbf{w}){\FFT}_n \mathbf{v}_{j} \nonumber\\
&= \langle \mathbf{v}^{\prime}_{i}, \mathbf{w}^{\prime}, \mathbf{v}^{\prime}_{j} \rangle \nonumber\\
&= \Re (\langle \mathbf{v}^{\prime}_{i}, \mathbf{w}^{\prime}, \mathbf{v}^{\prime}_{j} \rangle).
\end{align}
Here, $\mathbf{v}^{\prime}_{i}=\overline{{\FFT}_n\mathbf{v}_i}$, $\mathbf{v}^{\prime}_j={{\FFT}_n\mathbf{v}_j}$,
${\mathbf{w}^{\prime}}=\frac{1}{n}\diag({\FFT}_n\mathbf{w})$, and all of them are $n$ dimensional complex vectors. $\Re(\cdot)$ is the operator which takes the real parts of its argument. With this transformation, we can compute the bilinear transformation with a circulant matrix in $O(n \log n)$ using a FFT.

The DFT of an $n$-dimensional vector $\mathbf{x}$,  ${\FFT}_n\mathbf{x}$, is conjugate symmetric if and only if $\mathbf{x}$ is a real vector~\citep{hayashi:17-2}.
In our experiment, we initialize $\mathbf{w}'$ with the DFT of a real vector and update it in complex space.  We update ``frequency'' domain (complex space) vectors using only the operations which have correspondence to ``time domain'' vectors. Thus, as described in \cite{hayashi:17-2}, the conjugate symmetry of vectors are kept while learning because their initial values satisfy it.

\subsubsection{Expressiveness of the Bilinear Transformation Using Circulant Matrices}
In this section, we explain about the expressiveness of circulant matrices in relation to an arbitrary matrix $\mat{W}\in\Rset^{n\times n}$. \cite{complex-journal} show that for any $\mat{W}$, there exists a normal matrix $\mat{W}'\in\Cset^{n\times n}$ such that $\mat{W}=\Re(\mat{W}')$. Further, as with a symmetric matrix, a normal matrix can be diagonalized as follows:
\begin{equation*}
\mat{W}=\Re(\mat{W}')=\Re(\mat{O}\diag(\mat{w}')\mat{O}^{*}).
\end{equation*}
Here, $\mat{O}\in\Cset^{n\times n}$ is a unitary matrix, $\mat{O}^{*}$ is the conjugate transpose of $\mat{O}$ and $\mat{w}'\in\Cset^{n}$ is the complex vector which consists of the eigenvalues of $\mat{W}'$. The bilinear transformation whose weight matrix $\mat{W}$ is replaced with this, can be transformed into (\ref{eq:prop2}). The unitary matrix $\mat{O}$ is a bijective function, so the input vectors $\mat{v}_i$, $\mat{v}_j$ are learned as their one-to-one correspondent $\mathbf{v}^{\prime}_{i}=\mathbf{O}\transpose \mathbf{v}_i$, $\mathbf{v}^{\prime}_{j} = \mathbf{O}\transpose \mathbf{v}_j$, assuming that the unitary matrix $\mathbf{O}$ is learned implicitly.
Note that to simultaneously diagonalize the normal matrices whose real parts are the weight matrices $\mat{U}^{[1]},\mat{U}^{[2]},\dots,\mat{U}^{[L]}$ in the scoring functions for labels, we have to assume that they form a commuting family as with the discussion in \ref{sim-diag}.
\section{Related work}
\label{sect:related_works}
\subsection*{Dependency Parsing}

In recent years, various graph-based parsers with attention mechanisms have been proposed. 

\cite{mlp-attention} incorporated the attention mechanism used in machine translation~\citep{attention} into their graph-based parser. Their model receives vectors which are made by concatenating LSTM outputs corresponding to each word and its head candidates. Similarly, \cite{bilinear-attention} proposed a graph-based parser where they substitute the MLP-based classifier in~\citep{mlp-attention} with the bilinear one in their multi-task neural model, although they still use the MLP-based one in prediction of labels. Accordingly,~\cite{deepbiaffine} modified the model by~\cite{mlp-attention} using a biaffine classifier instead of an MLP-based one which enables the model to express not only the probability of a word receiving a particular word as dependent but also the prior probability of a word being a head. 

Likewise, in transition-based parsing literature, the state-of-the-art parser on the English Penn TreeBank by~\cite{stackpointer} uses an attention mechanism based on a biaffne classifier which calculates the probability distribution of the next word which comes into the stack at each time step, with LSTM outputs corresponding to each word in the input sentence. The models proposed in this paper can be incorporated into these models.

\subsection*{Parameter Reduction in Neural Networks}
Recently, numerous methods toward parameter reduction of neural networks have been proposed.

As a similar approach to proposed methods, there is a method where a projection matrix is decomposed into smaller matrices by lower-rank approximation~\citep{lu:16}. In addition, \cite{ishihara:18} introduced eigenvector decomposition into neural tensor networks ~\citep{ntn} and analyzed the effects of parameter reduction. Although the paper~\citep{ishihara:18} is similar to the present paper in that their methods address parameter reduction in the bilinear term, our work is different in that we apply it to deep biaffine parser.

There are some methods to reduce parameters in a projection matrix by sharing them. \cite{yu:15} used circulant matrices in the fully connected layers. Our models are different from theirs in that we use circulant matrices in the bilinear terms. ~\cite{chen:2015} perform parameter reduction with a hash kernel and ~\cite{vikas:15} with special matrices like toeplitz matrices. While these can be also used for bilinear terms, the methods based on real diagonalization and circulant matrices are superior to them in computational efficiency.

~\cite{hinton:15} proposed a model called distillation and were able to train a model which was more compact than the original one. However, it needs a lot of time for training since it needs to be trained again for distillation. ~\cite{itay:16} achieved a significant reduction of parameters by the quantization, but the reported accuracy is inferior to the original model. Theoretically, these methods can be combined with the proposed methods.

\section{Experiments}
\label{sec:experiments}

\begin{table*}[t!]
\centering
\begin{tabular}{lcccccc}
\toprule
\bf Treebank & \multicolumn{2}{c}{\bf Baseline} & \multicolumn{2}{c}{\bf Symmetry Matrix} &     \multicolumn{2}{c}{\bf Circulant Matrix} \\
\bf             & \bf UAS          & \bf LAS  & \bf UAS      & \bf LAS  & \bf UAS     & \bf LAS \\\cmidrule(lr){1-7}
UD\_Czech & 93.72 & 91.89 & 93.45(-0.27) & 91.50(-0.39) & \bf 93.87(+0.15) & \bf 92.01(+0.12) \\
UD\_German & 87.57 & 84.27 & 87.21(-0.36) & 83.91(-0.36) & \bf 87.61(+0.04) & \bf 84.39(+0.12) \\
UD\_English & \bf 91.05 & \bf 89.42 & 90.95(-0.1) & 89.22(-0.2) & 91.04(-0.01)& 89.31(-0.11) \\
UD\_Chinese & 87.67 & 85.58 & 87.55(-0.12) & 85.41(-0.17) & \bf 87.96(+0.29) & \bf 85.65(+0.07) \\
UD\_Slovenian-SST & 75.63 & 69.53 & 74.94(-0.69) & 68.58(-0.95) & \bf 75.90(+0.27) & \bf 70.24(+0.71) \\
UD\_Latin & 70.90 & 64.53 & 70.27(-0.63) & 63.18(-1.35) & \bf 72.38(+1.48) & \bf 66.00(+1.47)\\
UD\_French-ParTUT & 91.82 & 89.78 & \bf 92.09(+0.27) & 89.94(+0.16) & 91.94(+0.12) & \bf 90.09(+0.31) \\
UD\_Galician-TreeGal & 80.10 & 74.77 & 80.05(-0.05) & 74.78 (+0.01) & \bf 80.24(+0.14) & \bf 75.68(+0.91) \\

\bottomrule
\end{tabular}
\caption{Main results on CoNLL 2017 dataset.}
\label{results-ud}
\end{table*}

\begin{table*}[t!]
\centering
\begin{tabular}{lcccccc}
\toprule
\bf Reduced Samples & \multicolumn{2}{c}{\bf Baseline} & \multicolumn{2}{c}{\bf Symmetry Matrix} &     \multicolumn{2}{c}{\bf Circulant Matrix} \\
\bf             & \bf UAS          & \bf LAS  & \bf UAS      & \bf LAS  & \bf UAS     & \bf LAS \\\cmidrule(lr){1-7}
0 / 4 & \bf 91.05 & \bf 89.42 & 90.95(-0.1) & 89.22(-0.2) & 91.04(-0.01) & 89.31(-0.11) \\
1 / 4 & \bf 90.32 & \bf 88.57 & 90.05(-0.27) & 88.30(-0.27) & 90.29(-0.03) & 88.49(-0.08) \\
2 / 4 & 88.98 & 87.08 & \bf 89.15(+0.17) & \bf 87.17(+0.09) & 88.72(-0.26) & 86.63(-0.45) \\
3 / 4 & 87.24 & 85.08 & 87.27(+0.03) & 85.06(-0.02) & \bf 87.59(+0.35) & \bf 85.38(+0.3) \\
\bottomrule
\end{tabular}
\caption{Results in UD\_English with fewer training samples.}
\label{results-reduced}
\end{table*}

\begin{table*}[t!]
\centering
\begin{tabular}{lcccccc}
\toprule

\bf Treebank & \bf UAS & \bf LAS \\ \cmidrule(lr){1-3}

UD\_Slovenian-SST & 74.98(-0.65) & 69.07(-0.46) \\
UD\_Latin & 71.96(+1.06) & 65.68(+1.15) \\
UD\_French-ParTUT &  92.17(+0.35) & 90.44(+0.66) \\
UD\_Galician-TreeGal & 79.68(-0.42) & 74.82(+0.05) \\

\bottomrule
\end{tabular}
\caption{Baseline model with reduced dimensions.
The numbers in parentheses are the differences from the baseline model with full dimensions.}
\label{results-halfdim}
\end{table*}

\subsection{Dataset and Implementation}

We compared the models described above on several languages in the CoNLL 2017 shared task for Universal Dependency Parsing dataset. We chose four languages which have relatively abundant training examples: UD\_Chinese, UD\_Czech, UD\_English and UD\_German. And we also selected four languages which have fewer training examples: UD\_French-ParTUT, UD\_Galician-TreeGal, UD\_Latin and UD\_Slovenian-SST. 

As a baseline model, we used the dependency parser by Timothy Dozat\footnote{https://github.com/tdozat/Parser-v2} which achieved the highest accuracy on the shared task. The structures of the proposed models are based on the baseline model; we changed only the classifier part.

We only modified two hyper-parameters: we used no pretrained embeddings and initialized word embeddings with a uniform distribution. These settings remain the same throughout all experiments unless otherwise stated. 

We use gold word segmentation and gold POS tags while word segmentation and POS tagging are included in the shared task. We excluded these two tasks because the objective of this research is to show the effects of the proposed methods on biaffine classifiers which are not used for those tasks.

\subsection{Results}
We show the results of the baseline and two proposed methods in Table~\ref{results-ud}. The method based on circulant matrices outperformed the others on almost all languages except for English where the baseline model achieved the best accuracy and French-ParTUT where the method based on symmetric matrix did so in UAS. Interestingly, the method based on symmetric matrices underperformed the baseline on most languages. This might be because of the restricted expressiveness of a symmetric weight matrix in comparison to a circulant one especially in that the former is not appropriate for expressing asymmetric relations.

\begin{table}[t!] 
\centering
\begin{tabular}{lrrr}
\toprule
& \bf \shortstack{Number of \\ Parameters} & \bf Percentage\\ \cmidrule(lr){1-3}
Character LSTM & 241200 & 7.64\% \\
Bidirectional LSTM & 1927200 & 61.03\% \\
Arc MLP & 320800 & 10.16\% \\
Label MLP & 80200 & 2.54\% \\
Arc Classifier & 160400 & 5.08\% \\
Label Classifier & 377437 & 11.95\% \\
Others & 50400 & 1.60\% \\ \hline
TOTAL & 3157637 & 100\% \\
\bottomrule
\end{tabular}
\caption{\label{results} Percentage of baseline model parameters accounted for by each component.}
\label{table:ratio}
\end{table}

\begin{table*}[t!]
\centering
\begin{tabular}{lrrr}
\toprule
\bf  & \bf Baseline &  \bf Symmetry Matrix & \bf Circulant Matrix \\ \cmidrule(lr){1-4}
Arc Classifier & 160400& 1200& 1600\\
Label Classifier & 377437& 11100& 14800\\
Sum with shared parts & 3157637 & 2632100 & 2636200 \\ \cmidrule(lr){1-4}
Difference from the baseline & 0.0\%& \bf -16.64\%& \bf -16.51\%\\
\bottomrule
\end{tabular}
\caption{\label{results} Comparison of parameter sizes.}
 \label{table:comparison}
\end{table*}

\section{Analysis}

In this section, we examine the robustness to overfitting of the proposed methods.

\subsection{Relaxation of Overfitting}

First, to further examine how numbers of parameters affect the models, we conducted experiments on UD\_English treebank reducing the number of training examples by a quarter at a time.

The results of this experiment are shown in Table \ref{results-reduced}. Our methods performed better on smaller datasets where the number of training examples are less than or equal to a half of the original number of examples. This indicates that our methods not only became robust to overfitting through parameter reduction but also achieved high generalizability. 

Second, we ran the baseline model with classifier dimensions reduced from 400 to 200 for the arc classifier and 100 to 50 for the label classifier and compared it with the proposed methods on languages from small treebanks. As shown in Table \ref{results-halfdim}, simply reducing the number of dimensions hindered the accuracies in some languages while the method based on circulant matrix consistently outperformed the baseline model with the original number of dimensions in all of these languages from small treebanks as shown in Table \ref{results-halfdim}. This result indicates the effectiveness of the proposed method on smaller datasets.

\subsection{Parameter Reduction}
Table \ref{table:ratio} shows the proportion of the total parameters which each part of the baseline model account for. While LSTMs occupy the largest portion of the model, the second largest part is the classifiers which account for about 17\% of the parameters. Table \ref{table:comparison} indicates that the proposed methods were able to reduce the number of parameters by more than 16\%.

\subsection{Parsing Speed}
To test the parsing speed, we used an NVidia GTX1080 GPU and parsed the test dataset of UD\_English.  As mentioned in Section~\ref{sec:proposed}, both proposed methods are superior to the baseline model in terms of time complexity. Actually, the methods based on symmetry  took 13.86 seconds, circularity 15.06 and the baseline 17.76 seconds. These results are in accordance with theoretical time complexity.

\section{Conclusion}
In this paper, we reduced the number of parameters in the weight matrices in biaffine classifiers based on the assumption of symmetry or circularity and examined their effects on the CoNLL 2017 shared task for Universal Dependency Parsing dataset. As a result, the method based on circulant matrices outperformed the baseline model in most of languages with about 16\% parameter reduction. As future work, the L1 regularization method for CompLex \citep{Trouillon:2016:CES:3045390.3045609} proposed by \citep{DBLP:conf/aaai/ManabeHS18}  may be integrated into our methods to further reduce the number of parameters in the bilinear function. The script for the bilinear functions proposed here is provided on the GitHub page below\footnote{https://github.com/TomokiMatsuno/PACLIC32/blob/master/my\_linalg.py}.

\section{Acknowledgments} 
We are grateful to Michael Wentao Li for proofreading the present paper. We thank the anonymous reviewers. This work was supported by JSPS KAKENHI Grant Number JP18K11457 and JST CREST Grant Number JPMJCR1513, Japan.











\bibliographystyle{plainnat}
{\footnotesize
\bibliography{reference}
}

\end{document}